OpenStudy (anonymous):

Multiply or divide the rational expression, as indicated, and simplify your answer.

OpenStudy (anonymous):

\[\frac{z^2+2z+1}{2z^2+3z+1}* \frac{2z^2-5z-3}{z+1}\]

OpenStudy (hoblos):

x^2 +2x +1 = (x+1)^2 2x^2 +3x +1 = (2x+1)(x+1) 2x^2 -5x -3 = (2x+1)(x-3) replace them and simplify

OpenStudy (anonymous):

Wait what do i replace the one's to the right?

OpenStudy (anonymous):

where is my medal

OpenStudy (anonymous):

64

OpenStudy (anonymous):

Where did you get 64 lol?

OpenStudy (hoblos):

the final answer is x-3

OpenStudy (anonymous):

Thanks hoblos!